OpenStudy (anonymous):

Solve the Equation Round the answer to the nearest hundredth. 4.21x+5.39=12.07(2.01-4.72x)

OpenStudy (anonymous):

\[x -> 0.308 \]

OpenStudy (anonymous):

x=0.31

OpenStudy (anonymous):

Yes nzGeek, overlooked the rounding requirement.

OpenStudy (anonymous):

Show work please sorry forgot to add that

OpenStudy (anonymous):

5.39 + 4.21 x = 12.07 (2.01 - 4.72 x) 5.39 + 4.21 x = 24.2607 - 56.9704 x 56.9704 x + 4.21 x = 24.2607 - 5.39 x = 18.8707/61.1804 x = 0.308444 x = 0.31